# Spectral Manifold Harmonization for Graph Imbalanced Regression

### v1

@@ 6   Conclusion
6.1   Future Improvements
Spectral Manifold Harmonization (SMH) has shown strong potential for addressing imbalanced regression on graph-structured data, but several avenues remain for further enhancement. First, integrating domain-specific constraints into the graph generation process could improve the realism and scientific validity of the synthetic graphs. Second, the absence of semantic context in the current synthesis process limits the interpretability and relevance of the generated data, highlighting the need for a hybrid spectral-semantic approach. Future work will also involve evaluating SMH across a wider range of benchmark datasets and predictive models to further optimize performance. Additionally, we plan to conduct more comprehensive comparisons with existing state-of-the-art methods and expand the application domains beyond drug discovery, including areas such as biology and materials science, to better assess the generalizability of our method.
To this end, we aim to develop a hybrid framework that integrates semantic information into the generation and modeling pipeline to further enhance prediction performance and scientific relevance.
@@ 6.1   Future Improvements
Table 4: Hyperparameter search space
@@
Figure 4: Training and validation performance of property value prediction for each dataset for original and augmented training sets.

### v2

@@ 6. Conclusion
6.1. Future Improvements
Spectral Manifold Harmonization (SMH) has shown strong potential for addressing imbalanced regression on graph-structured data, but several avenues remain for further enhancement. First, integrating domain-specific constraints into the graph generation process could improve the realism and scientific validity of the synthetic graphs. Second, the absence of semantic context in the current synthesis process limits the interpretability and relevance of the generated data, highlighting the need for a hybrid spectral-semantic approach. Future work will also involve evaluating SMH across a wider range of benchmark datasets and predictive models to further optimize performance. Additionally, we plan to conduct more comprehensive comparisons with existing state-of-the-art methods and expand the application domains beyond drug discovery, including areas such as biology and materials science, to better assess the generalization of our method. To this end, we aim to develop a hybrid framework that integrates semantic information into the generation and modeling pipeline to further enhance prediction performance and scientific relevance.
@@ 6.1. Future Improvements
Table 3. Hyperparameter search space
@@
Figure 5. Training and validation performance of property value prediction for each dataset for original and augmented training sets.